in the number 160^4320 after how many digits from the right will you encounter the first odd digit?
Answer by satyanarayana(2)

solution:Hi.160^4320 can be written as (16^4320)*(10^4320).as 10^4320 will result in all zeroes.so last 4320 digits will be 0's.now 16^1=16.16^2=256,16^3=4096.Please note that last but one digit is odd for all 16 powers.so final result will be 4322th digit from right will be odd.since 4320 zeroes will be there at last.and correspondingly u will have some even number at last of 16 powers.The last but one digit will be odd for all powers of 16.So this is answer.
ANSWER IS 4322.